High End Crockery Corp set the following standards for their pottery bowls:

Std Qty or Hrs Std Rate or Price
Direct materials 5.5 lbs per bowl $6.50 per lb
Direct labor .4 hrs per bowl $22.50 per hour
Variable Overhead .4 hrs per bowl $2.50 per hour

The actual results ae as follows:

Actual output 8,250 units
Raw materials purchased 45,900 lbs
Raw materials used in production 44,750 lbs
Actual price raw materials $6.25 per lb
Actual direct labor hours 3,200 hours
Actual direct labor rate $22.75 per hour
Actual variable overhead rate $2.00

per hour

  1. What is the materials price variance?

  2. Is the materials price variance favorable or unfavorable? Indicate F or U as your answer

  1. What is the materials quantity variance?

  1. Is the materials quantity variance favorable or unfavorable? Indicate F or U as your answer

  1. What is the labor rate variance?

  2. Is the labor rate variance favorable or unfavorable? Indicate F or U as your answer

  3. What is the labor efficiency variance?

  4. Is the labor efficiency variance favorable or unfavorable? Indicate F or U as your answer

  5. What is the variable overhead rate variance?

  6. Is the variable overhead rate favorable or unfavorable? Indicate F or U as your answer

  7. What is the variable overhead efficiency variance?

  8. Is the variable overhead efficiency variance favorable or unfavorable? Indicate F or U as your answer
Answer #1

Workings:

Standard data for actual output:

Standard price = $6.50

Standard quantity = 8250*5.50 = 45375 lbs

Actual price = $6.25

Actual quantity = 44750 lbs

Standard Rate = $22.5

Actual Rate = $22.75

Standard hours = 8250*.4 = 3300 hours

Actual hours = 3200 hours

Standard variable overhead Rate = $2.50

Actual variable overhead Rate = $2.00

Solutions:

Materials price variance = (Standard price - Actual Price)*Actual quantity

=(6.50 - 6.25)*44750 = 11187.5 F

Material quantity variance = (Standard quantity - Actual quantity)*Standard price

=(45375-44750)*$6.5 = 4062.5 F

Labor rate variance = (Standard rate - Actual rate)*Actual hours

=(22.50 - 22.75)*3200 = 800 U

Labor efficiency variance = (Standard hours - Actual hours)*Standard rate

=(3300 - 3200)*$22.50 = $2250 F

Variable overhead rate variance = (Standard rate - Actual rate)*Actual hours

=(2.50 - 2.00)*3200 = 1600 F

Variable overhead efficiency variance = (Standard hours - Actual hours)*Standard rate

=(3300 - 3200)*$2.50 = 250 F